# Enhancing Target-unspecific Tasks through a Features Matrix

### v2

@@ Abstract
Recent developments in prompt learning of large vision-language models have significantly improved performance in target-specific tasks. However, these prompt optimizing methods often struggle to tackle the target-unspecific or generalizable tasks effectively. It may be attributed to the fact that overfitting training causes the model to forget its general knowledge having strong promotion on target-unspecific tasks. To alleviate this issue, we propose a novel Features Matrix (FM) regularization approach designed to enhance these models on target-unspecific tasks. Our method extracts and leverages general knowledge, shaping a Features Matrix (FM). Specifically, the FM captures the semantics of diverse inputs from a deep and fine perspective, preserving essential general knowledge, which mitigates the risk of overfitting. Representative evaluations demonstrate that: 1) the FM is compatible with existing frameworks as a generic and flexible module, and 2) the FM significantly showcases its effectiveness in enhancing target-unspecific tasks, achieving state-of-the-art performance.
@@
Figure 1: Our method is orthogonal to representative architectures, such as MaPLe (M) and PromptSRC (S), surpassing the existing state-of-the-art easy-to-use DePT by a significant margin.
Method Design Base Novel
CLIP (ICML2021) Hand 69.34 74.22
CoCoOp (CVPR2022) Tune 80.47 71.69
PLOT (ICLR2023) Tune 81.24 72.98
CoOp (IJCV2022) Tune 82.38 67.96
+ DePT (CVPR2024) Tune 83.66 71.82
+ Kg (CVPR2023) Tune + Hand 80.73 72.70
+ Ours Tune + Hand 81.15 74.66
@@ 1 Introduction
Large vision-language models such as CLIP  (Radford et al., 2021) have attracted increasing attention for remarkable generalization performance. Vision-language models (VLMs) are trained to align textual and visual modalities by leveraging extensive datasets. For instance, a prominent example of such models is CLIP, which has achieved remarkable success across a wide range of downstream tasks (Greer et al., 2024; Sanghi et al., 2021; Etchegaray et al., 2024). CLIP utilizes a large collection of 400 million text-image pairs. One of CLIP’s most appealing features is its ability to perform zero-shot inference. During inference, CLIP utilizes hand-crafted text inputs, known as prompts, to generate classification weights to predict image features, all without requiring any target-specific parameters training.
In contrast to hand-crafted prompts, a model-parameter tuning method has been proposed as prompt learning to automatically learn prompt embeddings (Zhou et al., 2022a). For instance, CoOp (Zhou et al., 2022b) represents the first method that specifically focuses on learning the text embeddings of prompts with few-shot samples training while keeping the CLIP model frozen. Although CoOp (Zhou et al., 2022b) has shown significant performance improvements over hand-crafted prompts in target-specific base classes, it may yield inferior performance compared to the hand-crafted prompt CLIP in generalization tasks, e.g., the novel class of generalization from base-to-novel (see Table 1).
@@
To overcome this challenge, some effective methods (Lu et al., 2022; Zhou et al., 2022a, b; Yao et al., 2023a; Chen et al., 2022) with tuning textual embeddings have been proposed. These methods aim to enhance the performance of novel classes, surpassing the novel performance of the previous CoOp. Although the novel ability of these methods surpasses that of CoOp ($67.96\%$), it is unfortunate that the novel ability of these methods still falls short compared to the hand-crafted prompt method (CLIP: $74.22\%$), as shown in Table 1 (Novel). A possible reason (Zhou et al., 2022a) may be that these prompting methods with tuning text embeddings tend to overfit the downstream data distributions. This overfitting can result in the model losing its inherent generalization capabilities (Yao et al., 2023a) obtained from hand-crafted prompts (Radford et al., 2021).
@@
To alleviate this problem, we propose a novel Features Matrix (FM) with CLIP for target-unspecific tasks. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained matrix from frozen CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts finely and deeply, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Importantly, our method is compatible with current prompt learning frameworks for textual or multi-modal prompt learning and serves as a flexible and generic module. As demonstrated in Figure 1, our method integrated to the representative Maple (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b), exhibits competitive results in base-to-novel generalization, domain generalization, and cross-dataset generalization across 11 datasets. Meanwhile, our method surpasses the existing state-of-the-art easy-to-use DePT (Zhang et al., 2024) by a significant margin. Our contributions can be summarized as follows:
@@
  • Various target-unspecific tasks (base-to-novel generalization, cross-dataset generalization, and domain generalization) across 11 datasets demonstrate that our method showcases its effectiveness.
Figure 2: Illustration of representative textual prompting frameworks (CoOp and CoCoOp) and multi-modal prompting frameworks (MaPLe and PromptSRC). We propose a flexible and generic design, which is compatible with these representative architectures. In the figure, ”snowflake pattern” represents parameter freezing, ”flame pattern” represents learnable pattern, ”Deep” represents learnable tokens embedded in several layers of the encoder, where ”T” represents learnable text embedding, ”V” represents learnable visual embedding, ”Class priors” represents which category it belongs to, ”Tuning Similarity” represents the calculation of cosine similarity for fine-tuning architecture, and the light gray ”Similarity” represents the calculation of cosine similarity for frozen architecture. In the MaPLe architecture diagram, the ”e” represents the matrix function that connects the encoders of two modalities in several layers.
@@ 2 Preliminaries
Textual Prompting. In textual prompting, the class name is retained as prior knowledge, while the word embeddings (referred to as context) of prompts are treated as learnable parameters, as shown in CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) of Figure 2. By modeling these context embeddings as trainable parameters, the model can optimize the prompts based on the specific requirements, enhancing the alignment between images and prompts. For class $k$, the tuning feature of the text encoder is denoted as ${t}^{tun}_{k}$ in a dataset with total $K$ classes. The final prediction probability of cross-entropy loss for two-modalities alignment is computed as:
@@
Multi-Modal Prompting. Based on textual deep prompting, this design uses only a limited number of trainable parameters in the image encoder, as shown in MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) of Figure 2. The visual prompts are introduced at every transformer layer’s input space. In the visual branch, the input image $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ is divided into $M$ patches. And, a tuning embedding of class $i_{class}$ is appended with the input patches. The foundation of vision prompting is ViT (Vision Transformer) (Dosovitskiy et al., 2020) backbone, which shares the same image encoder as CLIP. Let $\mathcal{P}_{v}$ denote the visual embeddings of prompts, and the image encoder processes the input tokens to generate tuning visual features, as follows (Jia et al., 2022):
$\tilde{\mathbf{x}}_{p}=\left\{\mathcal{P}_{v},\boldsymbol{i}_{class},\boldsymbol{i}_{1},\boldsymbol{i}_{2},\cdots,\boldsymbol{i}_{M}\right\},$ (3)
the tuning visual embeddings are introduced in the image encoder for deep prompt tuning.
@@ 3.1 Challenge and Motivation
As shown in Table 1, such as textual prompt (Zhou et al., 2022a, b; Yao et al., 2023a; Lu et al., 2022) and multi-modal prompting (Chen et al., 2022), have shown significant improvements in target-specific base classes. However, these methods often suffer from overfitting, leading to poor generalization in novel classes. To improve the novel performance of models, prompting regularization involves constraining the pre-trained and fine-tuning features in the text branch through computed loss. This constraint helps prevent forgetting general knowledge and ensures that the model retains essential information for downstream tasks. The objective of this design is to activate and maintain the pre-trained general knowledge, leveraging its remarkable generalization abilities to mitigate performance degradation when dealing with target-unspecific classes.
Recently, KgCoOp (Yao et al., 2023a) incorporates the constraint with $L1$ loss on fine-tuning feature with the pre-trained feature for the textual branch, avoiding the loss of CLIP’s original generalization capability. However, the novel performance of KgCoOp ($72.70\%$) is still lower than CLIP ($74.22\%$), as shown in Table 1 (Novel). It may be attributed to the fact that this single hand-crafted prompting regularization with CLIP cannot fully release and excavate the diversity semantics for utilizing essential general knowledge (Khattak et al., 2023b).
@@ 3.2 Proposed Features Matrix (FM)
To tackle this challenge and problem, we propose a novel design called Features Matrix (FM) for enhancing target-unspecific tasks. Our method incorporates multiple hand-crafted prompts with classes of the same datasets to extract general information as a pre-trained features matrix from CLIP to enhance generalization. In this way, we can explore and excavate the semantic information brought by each hand-crafted prompt more finely and deeply. By specifically aligning these multiple pre-trained hand-crafted unexpected features and tuning image features, our method aims to enhance the model’s generalization and robustness. This involves focusing on indistinguishable features that pose a challenge to the model during training.
As shown in Figure 3, we employ a set of hand-crafted prompt templates (P1, P2, P3, P4, etc) as inputs in the text encoder of pre-trained frozen CLIP. These prompt templates such as “a photo of one”, “a photo of a ”, “a picture of a”, “a drawing of a”, etc. These prompt templates are from the Prompt Engineering (Radford et al., 2021). In a dataset containing $K$ classes, the input to the text encoder is the frozen embeddings of hand-crafted prompts, consisting of different classes. The output of the frozen text encoder can be viewed as a features matrix, as depicted in Figure 3.
@@
Specifically, we let $t$ denote the sets of pre-trained text features extracted by a frozen text encoder through a set of hand-crafted prompts templates (P1, P2, P3, P4, etc) with classes. In a dataset with total $K$ classes, features belonging to the current label $k$ of image samples are referred to as designated features $t_{k}$, while features from other classes are considered non-designated features $t_{\hat{k}}$. Our method computes the matching scores with the cosine distance for each visual output feature ${v}^{tun}$ and designated pre-trained text features of the features matrix as $\operatorname{cos}({t}_{k},{v}^{tun})$. Similarly, we compute the matching scores between each visual output feature and the non-designated pre-trained text features, denoted as $\operatorname{cos}(t_{\hat{k}},{v}^{tun})$. In correspondence with the features matrix, these matching scores can be viewed as a scores matrix, as depicted in Figure 3. Subsequently, we identify the textual feature with lower scores (low-$\beta$) among the designated features, which is referred to as unexpected designated features set ${F}^{k}_{un}$. Similarly, we identify textual features with higher scores (top-$\beta$) among the non-designated features, which are referred to as unexpected non-designated feature sets ${F}^{\hat{k}}_{un}$. Accordingly, the contrastive loss $\mathcal{L}_{\text{CL}}$ of unexpected features similarity is denoted as follows,
$\mathcal{L}_{\text{CL}}=-\log\frac{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}}{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}+\exp\left\{\operatorname{cos}\left({t_{\hat{k}}},{v}^{tun}\right)\right\}},$ (4)
@@
where ${t}_{k}\in{F}^{k}_{un}$ and ${t_{\hat{k}}}\in{F}^{\hat{k}}_{un}$ as the selected unexpected features for designated unexpected features and non-designated unexpected features. The objective is to explore and excavate the unexpected features of general information. We train the objective for aligning the unexpected text features and tuning visual features, and the contrastive loss $\mathcal{L}_{\text{CL}}$ optimizes visual embeddings.
@@
Table 2: Base-to-novel generalization. Our method achieves consistent average performance improvement over different baselines.
@@
For two-modal alignment, let $\mathcal{L}_{\mathrm{CE}}(\cdot,\cdot)$ denote the cross-entropy loss with prompts $\mathbf{p}$ for samples $\mathcal{S}$, as follows,
Consequently, optimizing models with $\mathcal{L}_{\text{CL}}$ between the pre-trained textual unexpected features and image tuning features. The $\mathcal{L}_{\text{CE}}$ represents two modalities of alignment of deep vision-language prompting. Importantly, our method is compatible with representative prompt learning frameworks as a generic and flexible module for textual prompt learning, such as CoOp and CoCoOp, or multi-modal prompt learning such as MaPLe and PromptSRC.

### v4

@@ Abstract
Recent developments in prompt learning of large Vision-Language Models (VLMs) have significantly improved performance in target-specific tasks. However, these prompting methods often struggle to tackle the target-unspecific or generalizable tasks effectively. It may be attributed to the fact that overfitting training causes the model to forget its general knowledge. The general knowledge has a strong promotion on target-unspecific tasks. To alleviate this issue, we propose a novel Features Matrix (FM) approach designed to enhance these models on target-unspecific tasks. Our method extracts and leverages general knowledge, shaping a Features Matrix (FM). Specifically, the FM captures the semantics of diverse inputs from a deep and fine perspective, preserving essential general knowledge, which mitigates the risk of overfitting. Representative evaluations demonstrate that: 1) the FM is compatible with existing frameworks as a generic and flexible module, and 2) the FM significantly showcases its effectiveness in enhancing target-unspecific tasks (base-to-novel generalization, domain generalization, and cross-dataset generalization), achieving state-of-the-art performance.
@@ 1 Introduction
Large vision-language models such as CLIP  (Radford et al., 2021) have attracted increasing attention for remarkable generalization performance. Vision-Language Models (VLMs) are trained to align textual and visual modalities by leveraging extensive datasets. For instance, a prominent example of such models is CLIP, which has achieved remarkable success across a wide range of downstream tasks (Greer et al., 2024; Sanghi et al., 2022; Etchegaray et al., 2024). CLIP utilizes a large collection of 400 million text-image pairs. One of CLIP’s most appealing features is its ability to perform zero-shot inference. During inference, CLIP utilizes hand-crafted text inputs, known as prompts, to generate classification weights to predict image features, all without requiring any target-specific parameters training.
@@
In contrast to hand-crafted prompts, a model-parameter tuning method has been proposed as prompt learning to automatically learn prompt embeddings (Zhou et al., 2022a). For instance, CoOp (Zhou et al., 2022b) represents the first method that specifically focuses on learning the text embeddings of prompts with few-shot samples training while keeping the CLIP model frozen. Although CoOp (Zhou et al., 2022b) has shown significant performance improvements over hand-crafted prompts in target-specific base classes, it may yield inferior performance compared to the hand-crafted prompt CLIP in generalization tasks, e.g., the novel class of generalization from base-to-novel (see Table 1). To overcome this challenge, some effective methods (Lu et al., 2022; Zhu et al., 2023; Zhou et al., 2022a, b; Yao et al., 2023a; Chen et al., ; Li et al., 2024a) with tuning textual embeddings have been proposed. These methods aim to enhance the performance of novel classes, surpassing the novel performance of the previous CoOp. Although the novel ability of these methods surpasses that of CoOp ($67.96\%$), it is unfortunate that the novel ability of these methods still falls short compared to the hand-crafted prompt method (CLIP: $74.22\%$), as shown in Table 1 (Novel). A possible reason (Zhou et al., 2022a) may be that these methods with tuning text embeddings tend to overfit the downstream data distributions. This overfitting can result in the model losing its inherent generalization capabilities (Yao et al., 2023a) obtained from hand-crafted prompts (Radford et al., 2021).
@@
To alleviate this problem, we propose a novel Features Matrix (FM) with CLIP for target-unspecific tasks. Our method incorporates multiple hand-crafted prompts to extract general information as a pre-trained matrix from frozen CLIP to enhance generalization. The matrix of pre-trained features can delve into the semantics of different hand-crafted prompts finely and deeply, which reduces the risk of forgetting the essential general knowledge of pre-trained CLIP. Importantly, our method is compatible with current prompt learning frameworks for textual or multi-modal prompt learning and serves as a flexible and generic module. As demonstrated in Figure 1, our method, integrated with the representative MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b), exhibits competitive results in base-to-novel generalization, domain generalization, and cross-dataset generalization across 11 datasets. Meanwhile, our method surpasses the existing easy-to-use DePT (Zhang et al., 2024) by a significant margin. This underscores the robustness of our method across various tasks.
Our contributions can be summarized as follows:
Figure 1: Our method is orthogonal to representative architectures, such as MaPLe (M) and PromptSRC (S), surpassing the existing easy-to-use DePT (Zhang et al., 2024) by a significant margin.
@@
  • Various target-unspecific tasks (base-to-novel generalization, cross-dataset generalization, and domain generalization) across 11 datasets demonstrate that our method demonstrates its effectiveness.
Figure 2: Illustration of representative textual prompting frameworks (CoOp and CoCoOp) and multi-modal prompting frameworks (MaPLe and PromptSRC). We propose a flexible and generic design, which is compatible with these representative architectures. In the figure, “snowflake pattern” represents parameter freezing, “flame pattern” represents learnable pattern, “Deep” represents learnable tokens embedded in several layers of the encoder, where “T” represents learnable text embedding, “V” represents learnable visual embedding, “Class priors” represents which category it belongs to, “Tuning Similarity” represents the calculation of cosine similarity for fine-tuning architecture, and the light gray “Similarity” represents the calculation of cosine similarity for frozen architecture. In the MaPLe architecture diagram, the “e” represents the matrix function that connects the encoders of two modalities in several layers.
@@ 2 Preliminaries
Textual Prompting. In textual prompting, the class name is retained as prior knowledge, while the word embeddings (referred to as context) of prompts are treated as learnable parameters, as shown in CoOp (Zhou et al., 2022b) and CoCoOp (Zhou et al., 2022a) of Figure 2. By modeling these context embeddings as trainable parameters, the model can optimize the prompts based on the specific requirements, enhancing the alignment between images and prompts. For class $k$, the tuning feature of the text encoder is denoted as ${t}^{tun}_{k}$ in a dataset with a total of $K$ classes. The final prediction probability of cross-entropy loss for two-modalities alignment is computed as:
@@
Multi-Modal Prompting. Based on textual deep prompting, this design uses only a limited number of trainable parameters based on the image encoder, as shown in MaPLe (Khattak et al., 2023a) and PromptSRC (Khattak et al., 2023b) of Figure 2. The visual prompts are introduced at every transformer layer’s input space. In the visual branch, the input image $\mathbf{x}\in\mathbb{R}^{C\times H\times W}$ is divided into $M$ patches. And, a tuning embedding of class $i_{class}$ is appended with the input patches. The foundation of vision prompting is ViT (Vision Transformer) (Dosovitskiy et al., 2020) backbone, which shares the same image encoder as CLIP. Let $\mathcal{P}_{v}$ denote the visual embeddings of prompts, and the image encoder processes the input tokens to generate tuning visual features, as follows (Jia et al., 2022):
$\tilde{\mathbf{x}}_{p}=\left\{\mathcal{P}_{v},\boldsymbol{i}_{class},\boldsymbol{i}_{1},\boldsymbol{i}_{2},\cdots,\boldsymbol{i}_{M}\right\}.$ (3)
The tuning of visual embeddings are introduced in the image encoder for deep prompt tuning.
@@ 3.1 Challenge and Motivation
As shown in Table 1, such as textual prompting (Zhou et al., 2022a, b; Yao et al., 2023a; Lu et al., 2022) and multi-modal prompting (Chen et al., ; Li et al., 2024a), have shown significant improvements in target-specific base classes. However, these methods often suffer from overfitting, leading to poor generalization in novel classes. To improve the novel performance of models, prompting regularization involves constraining the pre-trained and fine-tuning features in the text branch through computed loss. This constraint helps prevent forgetting general knowledge and ensures that the model retains essential information for downstream tasks. The objective of this design is to activate and maintain the pre-trained general knowledge, leveraging its remarkable generalization abilities to mitigate performance degradation when dealing with target-unspecific classes.
Method Design Base Novel
CLIP (ICML2021) Hand 69.34 74.22
CoCoOp (CVPR2022) Tune 80.47 71.69
ProDA (CVPR2022) Tune 81.56 72.30
PLOT (ICLR2023) Tune 81.24 72.98
ProGrad (ICCV2023) Tune 82.48 70.75
WCA (ICML2024) Tune 84.30 74.01
CoOp (IJCV2022) Tune 82.38 67.96
+ DePT (CVPR2024) Tune 83.66 71.82
+ Ours Tune + Hand 81.15 74.66
It is important to recognize that the fixed CLIP’s fundamental pre-trained characteristics demonstrate robust generalization capabilities. Recently, KgCoOp (Yao et al., 2023a) incorporates the constraint with $L1$ loss on fine-tuning feature with the pre-trained feature for the textual branch, avoiding the loss of CLIP’s original generalization capability. However, the novel performance of KgCoOp ($72.70\%$) is still lower than pre-trained CLIP ($74.22\%$(Radford et al., 2021) of Table 1. It may be attributed to the fact that this single hand-crafted prompting regularization with CLIP cannot fully release and excavate the diverse semantics for utilizing essential general knowledge (Khattak et al., 2023b).
@@ 3.2 Proposed Features Matrix (FM)
To tackle this challenge and problem, we propose a novel design called Features Matrix (FM) for enhancing target-unspecific tasks. Our method incorporates multiple hand-crafted prompts with classes of the same datasets to extract general information as a pre-trained features matrix from CLIP to enhance generalization. In this way, we can explore and excavate the semantic information brought by each hand-crafted prompt more finely and deeply. By specifically aligning these multiple pre-trained hand-crafted unexpected features and tuning image features, our method aims to enhance the model’s generalization and robustness. This involves focusing on indistinguishable features that pose a challenge to the model during training. As shown in Figure 3, we employ a set of hand-crafted prompt templates (P1, P2, P3, P4, etc) as inputs in the text encoder of pre-trained frozen CLIP. These prompt templates, such as “a photo of one”, “a photo of a ”, “a picture of a”, “a drawing of a”, etc. The number of prompt templates (Radford et al., 2021) is set to 60. In a dataset containing $K$ classes, the input to the text encoder is the frozen embeddings of hand-crafted prompts, consisting of different classes. The output of the frozen text encoder can be viewed as a features matrix, as depicted in Figure 3. Specifically, we let $t$ denote the sets of pre-trained text features extracted by a frozen text encoder through a set of hand-crafted prompt templates (P1, P2, P3, P4, etc) with classes. In a dataset with a total of $K$ classes, features belonging to the current label $k$ of image samples are referred to as designated features $t_{k}$, while features from other classes are considered non-designated features $t_{\hat{k}}$.
@@
Our method computes the matching scores with the cosine distance for each visual output feature ${v}^{tun}$ and designated pre-trained text features of the features matrix as $\operatorname{cos}\ ({t}_{k},{v}^{tun})$. Similarly, we compute the matching scores between each visual output feature and the non-designated pre-trained text features, denoted as $\operatorname{cos}\ (t_{\hat{k}},{v}^{tun})$. In correspondence with the features matrix, these matching scores can be viewed as a scores matrix, as depicted in Figure 3. Subsequently, we identify the textual feature with lower scores (low-$\beta$) among the designated features, which is referred to as the unexpected designated features set ${F}^{k}_{un}$. The low-$\beta$ is that there is a list of scores sorted from low to high, with $\beta$ values assigned from the front of the list. Similarly, we identify textual features with higher scores (top-$\beta$) among the non-designated features, which are referred to as unexpected non-designated feature sets ${F}^{\hat{k}}_{un}$. Accordingly, the contrastive loss $\mathcal{L}_{\text{CL}}$ of unexpected features similarity is denoted as follows,
$-\log\frac{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}}{\exp\left\{\operatorname{cos}\left({t}_{k},{v}^{tun}\right)\right\}+\exp\left\{\operatorname{cos}\left({t_{\hat{k}}},{v}^{tun}\right)\right\}},$ (4)
@@
where the ${t}_{k}\in{F}^{k}_{un}$ and ${t_{\hat{k}}}\in{F}^{\hat{k}}_{un}$ denote the selected unexpected features for designated unexpected features and non-designated unexpected features. The objective is to explore and excavate the unexpected features of general information. We train the objective for aligning the unexpected text features and tuning visual features, and the contrastive loss $\mathcal{L}_{\text{CL}}$ optimizes visual embeddings. For two-modal alignment, let $\mathcal{L}_{\mathrm{CE}}(\cdot,\cdot)$ denote the cross-entropy loss with prompts $\mathbf{p}$ for samples $\mathcal{S}$, as follows,
Consequently, optimizing models with $\mathcal{L}_{\text{CL}}$ between the pre-trained textual unexpected features and image tuning features. The $\mathcal{L}_{\text{CE}}$ represents two modalities of alignment of deep vision-language prompting. Importantly, our method is compatible with representative prompt learning frameworks as a generic and flexible module for textual prompt learning, such as CoOp and CoCoOp, or multi-modal prompt learning, such as MaPLe and PromptSRC.
@@
Table 2: Base-to-novel generalization. Compared with FAP (NIPS2024) and an easy-to-use DePT (CVPR2024) in detail, our method achieves consistent average performance improvement over different representative baselines (CoOp, CoCoOp, MaPLe, PromptSRC).
@@
FAP 82.10 63.11 70.42